# Large Language Models as Neurolinguistic Subjects: Discrepancy in Performance and Competence for Form and Meaning

### v2

@@ Abstract
This study investigates the linguistic understanding of Large Language Models (LLMs) regarding signifier (form) and signified (meaning) by distinguishing two LLM assessment paradigms: psycholinguistic and neurolinguistic. Traditional psycholinguistic evaluations often reflect statistical rules that may not accurately represent LLMs’ true linguistic competence. We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pair and diagnostic probing to analyze activation patterns across model layers. This method allows for a detailed examination of how LLMs represent form and meaning, and whether these representations are consistent across languages. We found: (1) Psycholinguistic and neurolinguistic methods reveal that language performance and competence are distinct; (2) Direct probability measurement may not accurately assess linguistic competence; (3) Instruction tuning won’t change much competence but improve performance; (4) LLMs exhibit higher competence and performance in form compared to meaning. Additionally, we introduce new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets.
@@ 1 Introduction
Figure 1: Left: Research Question. Right: Psycholinguistic vs. Neurolinguistic Paradigm. Both direct probability measurement and metalinguistic prompting can be considered psycholinguistic methods, while minimal pair probing He et al. (2024) and other diagnostic probing are neurolinguistic.
There is currently a heated debate about whether LLMs understand human language or whether their performance is simply the product of complex statistical relationships Mitchell and Krakauer (2023). A central aspect of this debate concerns the nature of LLMs’ linguistic representations. Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how form and meaning intertwist with each other. Is LLM’s understanding of language meaning merely a statistical outcome based on their grasp of language form? When different languages express a shared concept with distinct forms, do LLMs create similar representations for these variations? How can we better understand the representations of form and meaning in these systems that support the observed patterns of performance?
@@ 2.2 In LLM Assessment Research
Computing a model’s probability of generating two minimally different sentences is one way to address these concerns Hu and Levy (2023). The minimal difference between the two sentences (e.g. replacement of a single word) makes one sentence acceptable while the other is not Linzen et al. (2016). Here are two examples for testing grammatical and conceptual understanding, respectively:
@@
Another line of research focuses on studying the internal representations, emphasizing a neurolinguistic approach to understanding LLMs. Essentially, diagnostic probing methods in evaluating language models can be considered as neurolinguistic paradigms as they examine the internal states of LMs Belinkov and Glass (2019); Belinkov (2022), while the term ‘neurolinguistic’ hasn’t been applied to the field before. Diagnostic probing involves training a classifier to predict linguistic properties from the hidden states of LMs. Following this paradigm, researchers decode syntactic, semantic, morphological, and other linguistic properties from the hidden states of LMs Köhn (2015); Gupta et al. (2015); Shi et al. (2016); Tenney et al. (2019); Manning et al. (2020).
@@ 3 Minimal Pair Probing = Minimal Pair + Diagnostic Probing
Specifically, given an LLM $f:x_{0,1,...,i}\rightarrow x_{i+1}$ trained on dataset $\mathcal{D}_{O}$, we can extract the hidden state representations $f_{l}(S)$ of the $l$-th layer of stimuli $S$. Given a minimal pair dataset $\mathcal{D}_{m}$ = $\{(S_{+}^{i},S_{-}^{i}),(z_{+}^{i},z_{-}^{i})\}$ with each sentence $S$ has a label $z$, we have internal representation $f_{l}(S_{+}^{i})$ and $f_{l}(S_{-}^{i})$ for each sentence. A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$. Note that our focus is the performance of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, instead of the capability of the supervised classifier $g$. To eliminate the influence of $g$, we set up a “random” baseline, where we assign a random vector $r$ to each sentence to replace the hidden states $f_{l}(S)$ and calculate the performance of the classifier $g$ on the random vector $r$ as $\mathrm{Perf}(g,\mathcal{D}_{m})$, aiming to evaluate the inherent predictability of the probing classifier. Finally, the performance of the LLM $f$ is defined as follows:
$\mathrm{Perf}(f,\mathcal{D}_{O})\triangleq\frac{\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})-\mathrm{Perf}(g,\mathcal{D}_{m})}{1-\mathrm{Perf}(g,\mathcal{D}_{m})}$ (1)
@@ 4.2 Setup for Psycholinguistic Analysis
Metalinguistic prompting involves explicitly asking a question or specifying a task that requires a judgement about a linguistic expression. Following Hu and Levy (2023), we use one prompt for a minimal pair to present both sentences at once. For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences and compare the probabilities assigned by the model to each answer option, “1” or “2”. For meaning tasks, we reformulate the property into a question and compare the probabilities of acceptable and unacceptable concepts as sentence continuations. Table 2 presents the prompts used in the experiments.
@@ 4.3 Setup for Neurolinguistic Analysis
We use logistic regression as the probing classifier and F1 score as the evaluation metrics. The score for $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$ and $\mathrm{Perf}(g,\mathcal{D}_{m})$ is calculated as the average F1 score across 5 cross-validation folds. Final performance $\mathrm{Perf}(f,\mathcal{D}_{O})$ is given by Formula 1.
@@ 5 Results
Figure 3: (a) Averaged psycholinguistic (meta and direct) and neurolinguistic results across models and tasks. t-tests were conducted on the original (pre-averaging) results between base and chat models, with p-values annotated. (b) Psycholinguistic and neurolinguistic performance for form (morphology, syntax-semantics interface, and syntax) and meaning (concept). (c) Competence and performance gap drops after instruction tuning.
@@ 5.1 Psycholinguistic vs Neurolinguistic
Figure 2 shows the performance of LLMs across all linguistic tasks. Figure 3-(a) demonstrates the averaged performance of LLMs across models and 4 levels (syntax, morphology, syntax-semantics interfaces, concept). Figure 3-(b) presents the average performance of LLMs across form and meaning tasks for Direct, Meta, and Neuro [^1] We refer to minimal pair probing as Neuro for simplicity. methods. We use the last layer’s performance in the Neuro method when comparing psycho- and neurolinguistic paradigms, as both direct probability measurement and metalinguistic prompting rely on the last layer of LLMs.
@@
Figure 3 shows distinct results between language performance and competence. Moving from Meta → Direct → Neuro, the evaluation focus gradually shifts from language performance (task execution ability) to language competence (the underlying linguistic ability). Within the same task category, Neuro methods consistently yield higher performance than Direct methods, which in turn outperform Meta methods. This indicates that when evaluating pure linguistic competence, LLMs perform well, but their performance drops when assessed in a task-based setting.
@@
For example, in the Neuro setting, performance on Syntax tasks reaches 97%, while in the Meta setting, it drops to 56.1%, showing a significant gap. This suggests that even when an LLM has strong competence in a given task, its performance can significantly decline when assessed under performance-oriented evaluation.
Direct probability measurement might not be true competence assessment.
As the Neuro method measures the internal representations of LLMs directly, it could serve as a reliable ground truth for estimating linguistic competence. Direct probability measurement falls short of achieving this ground truth in form assessment (especially for syntax and syntax-semantics-interface as shown in Figure 3-(b)).
@@
As shown in Figure 3-(b), LLMs consistently perform better on form-related tasks than on meaning-related tasks. This trend holds regardless of whether the model is a base or chat version. Crucially, this pattern is evident across all evaluation methods. This indicates that LLMs have a stronger grasp of linguistic form than conceptual meaning, whether assessed through task execution or underlying capability.
@@
Neuro results between the base and chat versions of LLMs reveal that instruction fine-tuning does not significantly alter the language competence of the models (t-test between Neuro-Base vs. Neuro-Chat as shown in Figure 3-(a)). With instruction fine-tuning (chat versions of LLMs), Meta performance on form improves significantly while meaning understanding remains stable. Figure 3-(c) illustrates that after instruction tuning, the competence-performance gap (Neuro-Meta) significantly decreases for form-related tasks, while the change for meaning-related tasks remains relatively small. This indicates that fine-tuning with well-designed instructions helps LLMs improve their performance on form-related tasks, bringing them closer to their underlying competence. However, for meaning-related tasks, instruction tuning does not lead to a fundamental improvement in understanding. This indicates that more optimized information access strategies can enhance the external performance of language models, particularly for form-related tasks.
5.2 Neurolinguistic Analysis [^2] Raw results for English, Chinese, and German can be found in Figure 8, 9 and 10 in the Appendix.
@@
The results in Figure 5-(c) show that the performance scores for conceptual understanding are significantly lower than those for grammatical understanding. This pattern is consistent across all six models, suggesting a universal characteristic of LLMs. Moreover, as illustrated in Figure 5-(a),(b), the encoding performance progressively declines from more structural tasks to more semantic tasks—spanning syntax, morphology, the syntax-semantic interface, and finally conceptual understanding. This highlights that LLMs encode features less effectively as the tasks shift from structure-focused to meaning-focused.
Figure 5: (a) Neuro probing results for 12 tasks in Llama-2 (4 syntax, 4 morphology, 4 syn.-sem. inter., 4 concept). (b) Averaged performance across four linguistic levels in Llama-2. (c) Averaged performance on form vs. meaning across all six models. (d) Feature learning saturation (95% of peak) and maximum layer. (e) Perf. difference between base and instruction-tuned LLMs. (f) Neuro probing results for Chinese and German. (g) Correlation between meaning competence and form competence.
@@
We compute the feature learning saturation and maximum layers for all 12 grammatical tasks and 4 conceptual tasks, averaging them to represent form and meaning, respectively. As shown in Figure 5-(d), the saturation and maximum layers for meaning are generally higher than those for form across all six models. This suggests that LLMs stabilize their encoding of grammatical features before conceptual features.
@@
As Figure 5-(e) shows, performance differences for form and meaning remain near zero across all layers, indicating that instruction tuning minimally impacts internal linguistic representations, consistent with our psycholinguistic vs. neurolinguistic analysis.
@@ 5.3 Multilingual analysis
How does LLMs’ understanding of meaning change when the form (language) varies? Since our COMPS-ZH and COMPS-DE datasets align with the concepts in the English COMPS dataset, we can explore whether LLMs’ grasp of different linguistic forms for the same concept correlates with their understanding of meaning across languages. Our previous results suggest that instruction tuning has little influence on the internal representations. Therefore we focus on the base LLMs here.
Cross-lingual Pattern
From Figure 5-(f), for all models and languages, form consistently achieves higher performance than meaning, indicating it’s easier for LLMs to make a stronger grasp of structural elements compared to conceptual comprehension. Extended multilingual results can be found in Appendix E.
Meaning competence is correlated to form competence.
Figure 5-(g) illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs. The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
@@ 6 Discussion
Language performance vs. competence: using neurolinguistic probing to reveal deeper competencies.
Our results demonstrate distinct patterns between the neurolinguistic and psycholinguistic evaluation methods. We observed that LLM performance evaluated using the Meta method was significantly lower than that of Direct, while Direct also performed notably worse than the Neuro method. This finding extends Hu and Levy (2023), who compared direct probability measurement (Direct) and meta-linguistic prompting (Meta) and argued that performance does not always reflect competence.
While Hu and Levy (2023) suggested that Direct probability measurement is a more reliable approach for assessing competence, our findings reveal that there is still a significant gap between Direct and Neuro probing, which directly examines internal representations. Hence, we recommend Neuro probing for evaluating competence.
Neuro probing further validates the discrepancy between performance and competence. While Meta results underperform, this does not necessarily indicate that the LLMs lack the underlying linguistic competence. Instead, it may reflect limitations in information access, as suboptimal prompts can prevent models from exhibiting their full capabilities. Specifically, prompting failures do not always equate to a lack of encoded knowledge. This observation is consistent with previous studies Firestone (2020); Lampinen (2024) that emphasize the importance of carefully controlling performance conditions for both humans and machines. methods can often reveal deeper competencies within the models.
@@
We further observe a linear correlation between form and meaning competence, particularly when linguistic forms vary across languages while meaning remains constant. This suggests that LLMs’ understanding of meaning might rely heavily on form, with conceptual representation anchored to formal structures rather than independent meaning comprehension. These results offer a semiotic and neurolinguistic explanation for LLMs’ long-standing issue of generating “confidently incorrect” responses, i.e. hallucinations Ji et al. (2023).

### v3

# Large Language Models as Neurolinguistic Subjects: Discrepancy between Performance and Competence

@@ Abstract
This study investigates the linguistic understanding of Large Language Models (LLMs) regarding signifier (form) and signified (meaning) by distinguishing two LLM assessment paradigms: psycholinguistic and neurolinguistic. Traditional psycholinguistic evaluations often reflect statistical rules that may not accurately represent LLMs’ true linguistic competence. We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pairs and diagnostic probing to analyze activation patterns across model layers. This method allows for a detailed examination of how LLMs represent form and meaning, and whether these representations are consistent across languages. We found: (1) Psycholinguistic and neurolinguistic methods reveal that language performance and competence are distinct; (2) Direct probability measurement may not accurately assess linguistic competence; (3) Instruction tuning won’t change much competence but improve performance; (4) LLMs exhibit higher competence and performance in form compared to meaning. Additionally, we introduce new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets. [^1] Code and data available here.
@@ 1 Introduction
Figure 1: Illustration of LLMs processing the same signified (meaning) across different signifiers (forms).
Figure 2: Psycholinguistic vs. Neurolinguistic Paradigm. Both direct probability measurement and metalinguistic prompting can be considered as psycholinguistic methods, while minimal pair probing He et al. (2024) and other diagnostic probing are neurolinguistic.
There is currently a heated debate about whether LLMs understand human language or whether their performance is simply the product of complex statistical relationships Mitchell and Krakauer (2023). A central aspect of this debate concerns the nature of LLMs’ linguistic representations. Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how form and meaning intertwist with each other. Is LLMs‘ understanding of language meaning merely a statistical outcome based on their grasp of language form? When different languages express a shared concept with distinct forms, do LLMs create similar representations for these variations? How can we better understand the representations of form and meaning in these systems that support the observed patterns of performance?
@@ 2.2 In LLM Assessment Research
Computing a model’s probability of generating two minimally different sentences is one way to address these concerns Hu and Levy (2023). The minimal difference between the two sentences (e.g., replacement of a single word) makes one sentence acceptable while the other is not Linzen et al. (2016). Here are two examples for testing grammatical and conceptual understanding, respectively:
@@
Another line of research focuses on studying the internal representations, emphasizing a neurolinguistic approach to understanding LLMs. Essentially, diagnostic probing methods in evaluating language models can be considered as neurolinguistic paradigms as they examine the internal states of LMs Belinkov and Glass (2019); Belinkov (2022), while the term ‘neurolinguistic’ hasn’t been applied to the field before. Diagnostic probing involves training a classifier to predict linguistic properties from the hidden states of LMs. Following this paradigm, researchers decode syntactic, semantic, morphological, and other linguistic properties from the hidden states of LMs Köhn (2015); Gupta et al. (2015); Shi et al. (2016); Tenney et al. (2019); Hewitt and Manning (2019); Manning et al. (2020).
@@ 3 Minimal Pair Probing = Minimal Pair + Diagnostic Probing
Specifically, given an LLM $f:x_{0,1,...,i}\rightarrow x_{i+1}$ trained on dataset $\mathcal{D}_{O}$, we can extract the hidden state representations $f_{l}(S)$ of the $l$-th layer of stimuli $S$. Given a minimal pair dataset $\mathcal{D}_{m}$ = $\{(S_{+}^{i},S_{-}^{i}),(z_{+}^{i},z_{-}^{i})\}$ with each sentence $S$ has a label $z$, we have internal representation $f_{l}(S_{+}^{i})$ and $f_{l}(S_{-}^{i})$ for each sentence. A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$.
Note that our focus is on evaluating the linguistic competence of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, rather than the capacity of the probing classifier $g$. As suggested by Hewitt and Liang (2019), even untrained or random representations can yield surprisingly high probing accuracy, raising concerns that the classifier may exploit dataset artifacts rather than meaningful representations. To control for the potential bias introduced by $g$, we construct a random embedding baseline.
Specifically, for each sentence in the dataset, we assign a fixed random vector $r$, sampled from a Gaussian distribution with the same mean and standard deviation as the real model embeddings $f_{l}(S)$. Importantly, each sentence is consistently assigned the same random vector across occurrences, preserving instance-level identity that the probing classifier might exploit. This allows us to assess the extent to which task performance can be driven by superficial sentence-level cues rather than meaningful representations. We then compute $\mathrm{Perf}(g,\mathcal{D}_{m})$ by training $g$ on these random embeddings, which reflects the inherent predictability or “shortcut” potential of the probing task. Therefore, our performance score incorporates a correction factor based on this random baseline, defined as:
$\mathrm{Perf}(f,\mathcal{D}_{O})\triangleq\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})\cdot(1+\frac{0.5-\mathrm{Perf}(g,\mathcal{D}_{m})}{0.5})$ (1)
This formula applies a correction term, penalizing cases where the probing classifier performs well even on random embeddings. When $\mathrm{Perf}(g,\mathcal{D}_{m})=0.5$, the correction factor is 1; if the performance is higher, the factor shrinks toward 0, discouraging overfitting or trivial tasks; if it drops below 0.5, the factor exceeds 1, slightly amplifying the model’s score. This ensures that only meaningful representations in $f$ contribute to the final evaluation.
@@ 4.2 Setup for Psycholinguistic Analysis
Metalinguistic prompting involves explicitly asking a question or specifying a task that requires a judgement about a linguistic expression. Following Hu and Levy (2023), we use one prompt for a minimal pair to present both sentences at once. For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences, and compare the probabilities assigned by the model to each answer option, “1” or “2”. For meaning tasks, we reformulate the property into a question and compare the probabilities of acceptable and unacceptable concepts as sentence continuations. Table 2 presents the prompts used in the experiments.
@@ 4.3 Setup for Neurolinguistic Analysis
We use logistic regression as the probing classifier and F1 score as the evaluation metric. The score for $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$ and $\mathrm{Perf}(g,\mathcal{D}_{m})$ is calculated as the average F1 score across 5 cross-validation folds. Final performance $\mathrm{Perf}(f,\mathcal{D}_{O})$ is given by Formula 1.
@@ 5.1 Psycholinguistic vs Neurolinguistic
Figure 3 shows the performance of LLMs across all linguistic tasks. Figure 4 demonstrates the averaged performance of LLMs across models and 4 levels (syntax, morphology, syntax-semantics interfaces, concept). Figure 5 presents the average performance of LLMs across form and meaning tasks for Direct, Meta, and Neuro [^2] We refer to minimal pair probing as Neuro for simplicity. methods. We use the last layer’s performance in the Neuro method when comparing psycho- and neurolinguistic paradigms, as both direct probability measurement and metalinguistic prompting rely on the last layer of LLMs.
@@
Figure 4 and 5 shows distinct results between language performance and competence. Moving from Meta → Direct → Neuro, the evaluation focus gradually shifts from language performance (task execution ability) to language competence (the underlying linguistic ability). Within the same task category, Neuro methods consistently yield higher performance than Direct methods, which in turn outperform Meta methods. This indicates that when evaluating pure linguistic competence, LLMs perform well, but their performance drops when assessed in a task-based setting.
Figure 4: Averaged psycholinguistic (meta and direct) and neurolinguistic results across models and tasks. t-tests were conducted on the original (pre-averaging) results between base and chat models, with p-values annotated.
@@
For example, in the Neuro setting, performance on Syntax tasks reaches 97%, while in the Meta setting, it drops to 56.1%, showing a significant gap. This suggests that even when an LLM has strong competence in a given task, its performance can significantly decline when assessed under a performance-oriented evaluation.
Figure 5: Psycholinguistic and neurolinguistic performance for form (morphology, syntax-semantics interface, and syntax) and meaning (concept).
Direct probability measurement might not be a true competence assessment.
As the Neuro method measures the internal representations of LLMs directly, it could serve as a reliable ground truth for estimating linguistic competence. Direct probability measurement falls short of achieving this ground truth in form assessment (especially for syntax and syntax-semantics-interface as shown in Figure 5).
Figure 6: Competence and performance gap drops after instruction tuning.
@@
As shown in Figure 5, LLMs consistently perform better on form-related tasks than on meaning-related tasks. This trend holds regardless of whether the model is a base or chat version. Crucially, this pattern is evident across all evaluation methods. This indicates that LLMs have a stronger grasp of linguistic form than conceptual meaning, whether assessed through task execution or underlying capability.
@@
Neuro results between the base and chat versions of LLMs reveal that instruction fine-tuning does not significantly alter the language competence of the models (t-test between Neuro-Base vs. Neuro-Chat as shown in Figure 4). With instruction fine-tuning (chat versions of LLMs), Meta performance on form improves significantly while meaning understanding remains stable. Figure 6 illustrates that after instruction tuning, the competence-performance gap (Neuro-Meta) significantly decreases for form-related tasks, while the change for meaning-related tasks remains relatively small. This indicates that fine-tuning with well-designed instructions helps LLMs improve their performance on form-related tasks, bringing them closer to their underlying competence. However, for meaning-related tasks, instruction tuning does not lead to a fundamental improvement in understanding. This indicates that more optimized information access strategies can enhance the external performance of language models, particularly for form-related tasks.
Figure 8: (a) Neurolinguistic probing performance for 16 tasks in Llama-2, including 4 syntax tasks, 4 morphology tasks, 4 syntax-semantics interface tasks, and 4 conceptual tasks. (b) Average probing performance across the four linguistic categories in Llama-2. (c) Mean performance comparison between form-related tasks (syntax, morphology, syntax-semantics interface) and meaning-related tasks (concept), aggregated across all six models.
Figure 9: Feature learning saturation layer (defined as the first layer reaching 95% of peak performance) and the layer of maximum performance.
Figure 10: Difference in probing performance between base and instruction-tuned models across all layers.
5.2 Neurolinguistic Analysis [^3] Raw results for English, Chinese, and German can be found in Figure 15, 16 and 17 in the Appendix.
@@
The results in Figure 8-(c) show that the performance scores for conceptual understanding are significantly lower than those for grammatical understanding. This pattern is consistent across all six models, suggesting a universal characteristic of LLMs. Moreover, as illustrated in Figure 8-(a),(b), the encoding performance progressively declines from more structural tasks to more semantic tasks, spanning syntax, morphology, the syntax-semantic interface, and finally conceptual understanding. This highlights that LLMs encode features less effectively as the tasks shift from structure-focused to meaning-focused.
@@
We compute the feature learning saturation and maximum layers for all 12 grammatical tasks and 4 conceptual tasks, averaging them to represent form and meaning, respectively. As shown in Figure 9, the saturation and maximum layers for meaning are generally higher than those for form across all six models. This suggests that LLMs stabilize their encoding of grammatical features before conceptual features.
@@
As Figure 10 shows, performance differences for form and meaning remain near zero across all layers, indicating that instruction tuning minimally impacts internal linguistic representations, consistent with our psycholinguistic vs. neurolinguistic analysis.
@@ 5.3 Multilingual analysis
Figure 11: Neuro probing results for English, Chinese, and German.
How does LLMs’ understanding of meaning change when the form (language) varies? Since our COMPS-ZH and COMPS-DE datasets align with the concepts in the English COMPS dataset, we can explore whether LLMs’ grasp of different linguistic forms for the same concept correlates with their understanding of meaning across languages. Our previous results suggest that instruction tuning has little influence on the internal representations. Therefore, we focus on the base LLMs here.
From Figure 11, for all models and languages, form consistently achieves higher performance than meaning, indicating it’s easier for LLMs to make a stronger grasp of structural elements compared to conceptual comprehension. Extended multilingual analysis can be found in Appendix E.
@@ 6 Discussion
Language performance vs. competence: probing reveals deeper linguistic understanding than direct probability.
Our results demonstrate that neurolinguistic probing uncovers linguistic competencies in LLMs that are not captured by psycholinguistic methods. While Meta performs the worst and Direct performs better, Neuro consistently outperforms both, revealing a systematic underestimation of competence when relying on output-based evaluations.
Hu and Levy (2023) argued that Direct probability measurement, being more intrinsic than metalinguistic prompting, better reflects competence. However, our findings show that even Direct falls short of revealing the full extent of LLMs’ linguistic capabilities. Direct relies on the final output layer, which is highly optimized for next-word prediction and thus entangled with task-specific objectives. Prior studies Hewitt and Manning (2019); Liu et al. (2019) have shown that syntactic and general linguistic information is often better represented in intermediate layers than in the final layer. Waldis et al. (2024) also emphasized that output correctness is an insufficient indicator of linguistic understanding, advocating for probing internal representations.
Our t-SNE visualizations corroborate this: clear linguistic clusters emerge in intermediate layers but dissolve in the final layer, reinforcing the view that the last layer is not optimal for assessing competence. These findings suggest that Direct, while more grounded than prompting, is still a limited proxy for internal knowledge.
In contrast, neurolinguistic probing inspects internal activation patterns across layers and tasks, uncovering the underlying representational structure of form and meaning, and further validates the discrepancy between performance and competence.
On the other hand, while Meta results underperform, this does not necessarily indicate that the LLMs lack the underlying linguistic competence. Instead, it may reflect limitations in information access, as suboptimal prompts can prevent models from exhibiting their full capabilities. Specifically, prompting failures do not always equate to a lack of encoded knowledge. This aligns with prior work Firestone (2020); Lampinen (2024) emphasizing the need to distinguish performance conditions from underlying ability.
Thus, we argue that probing, particularly when applied layer-wise, provides a more accurate and comprehensive assessment of linguistic competence than Direct probability alone.
@@
We further observe a linear correlation between form and meaning competence, particularly when linguistic forms vary across languages while meaning remains constant. This suggests that LLMs’ understanding of meaning might rely heavily on form, with conceptual representation anchored to formal structures rather than independent meaning comprehension. These results offer a semiotic and neurolinguistic explanation for LLMs’ long-standing issue of generating “confidently incorrect” responses, i.e., hallucinations Ji et al. (2023).